How many electrons are present in 180gm of water. Please help me out fast

Grade:11

1 Answers

Priyam raj
6 years ago
Hello mr future iitian.....In one molecule of water there are total 1+1+8=10 electrons..and in mole of water there are Na * 10 electrons...so now we have 180 gm of water....so first calculate no of moles i.e 180/(18) = 10 moles...Therefore required no of electrons= 10 * Na * 10= 10*6*10^23*10 electrons(calculate this on your copy)!Thank u too much ...